# The Influence of Clouds and Deuterium-Burning on Brown Dwarf Habitable Zones

### v1

@@ III.3 The Effects of Metallicity on Brown Dwarf Habitable Zones
III.4 Inner and Outer Habitable Zone Limits
We visualize the trajectory of the inner and outer edges of brown dwarf HZs (both cloudless and cloudy evolution) through time in Figure 9. The more massive the brown dwarf, the further-out the HZ is at 0.001 Gyr compared to less massive brown dwarfs. It also shows that the HZ shrinks as the brown dwarf ages. For example, for cloudless 0.012, 0.050, and 0.080 M⊙ brown dwarfs, the HZ starts off with a width of 7.06 AU, 1.70 AU, and 2.77 AU, respectively, at 0.001 Gyr. At 0.1 Gyr, the width of the HZ for those same brown dwarfs are 2.64 AU, 3.13 AU, and 5.20 AU, respectively. The figure demonstrates that the habitable zone location is highly time-dependent, shifting in radius substantially over the brown dwarf’s lifetime.
The corotation radius is the orbital distance where a planet’s orbital period matches the rotation period of its host. Planets inside this radius orbit faster than the host’s rotation, leading to tidal torques that drive inward migration (Bolmont et al. 2011). Conversely, planets outside the corotation radius orbit more slowly and experience outward tidal migration. Figure 9 (gray shaded regio000000000000000n) shows the corotation radii for brown dwarfs of 0.012, 0.050, and 0.080 M⊙, computed from rotation periods ranging from 5–15 hours. For the M dwarf (0.080 M⊙), the HZ remains entirely beyond the corotation radius at all ages.
However, for the 0.012 and 0.050 M⊙ brown dwarfs, the habitable zone migrates inward over time and intersects the corotation radius from Gyr (Figure 9). Any planet residing within what will become the habitable zone are expected to have already experienced strong tidal torques that drive inward orbital decay. As a result, planets may be removed through in spiral or disruption before the habitable zone completes its inward evolution, potentially leaving no planets remaining within the habitable zone at later times. Future studies must integrate both brown dwarf evolution and tidal migration with the modeling framework utilized in this work. This raises the question of which planets could actually have a lifetime in the habitable zone while residing within the corotation radius. Table 1 shows the calculated corotation radii for a select few brown dwarfs to provide reference. The corotation radius was calculated using Equation 3.
@@ IV Discussion
As shown in section III.1, previous studies such as Lingam et al. (2020) relied on analytic approximations rather than detailed brown dwarf evolution models, which did not accurately illustrate substellar HZ characteristics. Figure 1 illustrates this difference: at early ages, brown dwarfs are considerably cooler than the analytic models predict and thus much less likely to lose their entire water inventory than in the analytic case, although atmospheric escape timescales from brown dwarf planets requires further study. For the first time, we have computed equilibrium tracks for planets orbiting brown dwarfs using realistic, modern models that provide brown dwarf radii and luminosity tracks starting at 1 Myr. This permits us to explore the impact of clouds and deuterium burning on brown dwarf HZs. Since clouds slow the evolution of brown dwarfs, the cloudy evolution tracks result in orbiting planets that cool more slowly over time.
@@
Moreover, the transition from cloudy to cloudless states in the brown dwarf further delays planetary cooling at early ages, allowing well placed planets to maintain equilibrium temperatures in the HZ range for longer than predicted for cloudless evolution models. Lingam et al. (2020) concluded that planets orbiting brown dwarfs with masses less than 0.030 M⊙ would not maintain habitable temperatures over significant geologic timescales. This work illustrates that if cloud formation/dissipation, deuterium burning, and realistic atmospheric opacities are incorporated into atmospheric models, planets could remain in the habitable zone for hundreds of millions to billions of years (depending on the brown dwarf physical parameters, orbital radius, etc.), including planets orbiting brown dwarfs less than 0.030 M⊙ (shown in Figures 8 and 6).
@@
Sections III.2 and III.3 examined the duration that planets can remain within the HZ as a function of brown dwarf properties and orbital distance. Since higher-metallicity brown dwarfs cool more slowly and remain warmer at older ages than their lower-metallicity counterparts (e.g., Marley et al., 2021), planets orbiting metal-rich brown dwarfs spend longer periods within the HZ (see Figures 5 and 7). Consequently, planets orbiting metal-rich brown dwarfs spend longer periods within the HZ. This trend is evident in Figures 8 and 6, where systems with [M/H] = +0.5 exhibit extended habitable lifetimes compared to those with lower metallicities. The HZ lifetime also depends strongly on orbital distance and host mass: planets on wider orbits generally maintain habitability longer, particularly around lower-mass brown dwarfs. Figure 6 also shows the impact clouds have habitable zone lifetime. In the 0.001 AU case shown in Figures 8 and 6, planets with masses above approximately 0.050 M⊙ exhibit longer lifetimes in Figure 8 than in Figure 6. This difference arises because the brown dwarfs at these masses are hotter (due to the formation of clouds), resulting in planetary temperatures exceeding 275 K.
@@
Section III.4 examines the evolution of the inner and outer habitable zone boundaries for both cloudless and cloudy brown dwarf models. In all cases, the habitable zone contracts and migrates inward as the brown dwarf cools over time. The inward migration and contraction of brown dwarf habitable zones shown in Figure 9 have important consequences when coupled with tidal evolution. For the 0.012 and 0.050 M brown dwarfs, the habitable zone intersects the corotation radius at late times, implying that planets residing in the habitable zone may experience strong inward tidal migration and be removed before long-term habitability can be sustained. In contrast, for the 0.080 M M dwarf, the habitable zone remains exterior to the corotation radius throughout its evolution.
@@
It is also important to note that connecting the hydrogen-burning limit of main-sequence stars to the brown dwarfs shifts the brown dwarf habitable zone farther outward, meaning the corresponding temperatures are higher. This occurs because Kopparapu et al. (2013) adopt lower albedos (and used a full radiative-convective model), which trigger a runaway greenhouse at lower levels of insolation and therefore place the inner edge of the habitable zone at smaller orbital distances. The planetary albedo vs. surface temperature of planets orbiting a variety of main-sequence stars is shown in Figure 6 in Kopparapu et al. (2013).
@@
In contrast, the shallower luminosity–mass relation of brown dwarfs produces nearly parallel HZ boundaries, yielding a relatively constant absolute width at younger ages. The systematically larger orbital distances of the main-sequence HZs are a direct consequence of their higher effective temperatures and bolometric luminosities, which shift the equilibrium temperature range for liquid water to wider separations. It is also clear that the HZ boundaries shift inwards with time since brown dwarfs cool with age. Overall, brown dwarf HZs sweep through a large range of radii and Figure 10 illustrates how complex the evolution is.
@@
Although no rocky exoplanets have yet been discovered around brown dwarfs, studies have placed theoretical constraints on the occurrence rates of short-period planets in such systems (e.g., He et al. (2017)). The transit method is particularly favorable for detecting rocky planets orbiting brown dwarfs: an Earth-sized planet transiting a brown dwarf with a radius of produces a transit depth of ppm, compared to only ppm for the same planet orbiting a G-type star. This large contrast underscores the strong potential for future searches targeting habitable-zone planets around brown dwarfs. In future work, we will examine how the unique spectral energy distributions of brown dwarfs shape planetary climate by applying a 1D radiative–convective model to determine the inner (runaway greenhouse) and outer habitable zone limits for Earth-like planets (Windsor et al., 2022).
@@
However, the variability of the brown dwarf can cause the transit data to be extremely noisy. This is also seen in M-dwarf variability. There have been many studies analyzing the high activity and its effect on transits in M-dwarfs (e.g., Yaptangco et al., 2025; Miyakawa et al., 2022). Similar studies to those listed could be a point of research for brown dwarfs as well. Considering that brown dwarfs have opposite evolution as main-sequence stars could provide another interesting window to the possibilities of habitability beyond Earth.
@@
While (e.g., Barnes and Heller, 2013) highlighted fundamental challenges to habitability around cooling brown dwarfs (including tides), our work demonstrates that when realistic brown dwarf physics, deuterium burning, cloud evolution, and metallicity-dependent cooling, is included, habitable zone lifetimes can be substantially longer than previously assumed. In contrast to (e.g., Bolmont, 2018), which emphasized that inward habitable zone migration coupled with tidal torques near the corotation radius may rapidly eliminate potentially habitable planets around brown dwarfs, our results demonstrate that when realistic brown dwarf evolutionary physics is taken into account, planets located beyond AU, depending on the mass and corotation radius of the host brown dwarf, can both remain outside the corotation radius and persist within the habitable zone for extended durations (shown in Table 1).
@@ V Conclusions
We find that brown dwarfs are significantly less luminous at young ages than predicted by analytic approximations used in some recent brown dwarf habitable zone studies (Lingam et al., 2020). Those approximations neglect deuterium burning and cloud formation and dissipation, both of which play critical roles in determining accurate habitable zone boundaries and their evolution.
Deuterium burning and cloud formation and dissipation substantially extend habitable zone lifetimes for planets at specific orbital distances by maintaining those planets within the habitable zone while these processes operate.
Brown dwarfs with higher metallicities support longer habitable zone lifetimes because increased atmospheric opacity slows their cooling.
As brown dwarfs age, their habitable zones contract by orders of magnitude, depending on the brown dwarf mass.

### v2

@@ III.3 The Effects of Metallicity on Brown Dwarf HZs
III.4 Inner and Outer HZ Limits
We visualize the trajectory of the inner and outer edges of brown dwarf HZs (both cloudless and cloudy evolution) through time in Figure 9. The more massive the brown dwarf, the further-out the HZ is at 0.001 Gyr compared to less massive brown dwarfs. It also shows that the HZ shrinks as the brown dwarf ages. For example, for cloudless 0.012, 0.050, and 0.080 M⊙ brown dwarfs, the HZ starts off with a width of 7.06 au, 1.70 au, and 2.77 AU, respectively, at 0.001 Gyr. At 0.1 Gyr, the width of the HZ for those same brown dwarfs are 2.64 au, 3.13 au, and 5.20 au, respectively. The figure demonstrates that the HZ location is highly time-dependent, shifting in radius substantially over the brown dwarf’s lifetime. This change in width is also prominent in the cloudy evolution case, particularly for 0.012 and 0.050 M⊙.
The corotation radius is the orbital distance where a planet’s orbital period matches the rotation period of its host. Planets inside this radius orbit faster than the host’s rotation, leading to tidal torques that drive inward migration (Bolmont et al. 2011). Conversely, planets outside the corotation radius orbit more slowly and experience outward tidal migration. Figure 9 (gray shaded region) shows the corotation radii for brown dwarfs of 0.012, 0.050, and 0.080 M⊙, computed from rotation periods ranging from 5– to 15 hr. For the M dwarf (0.080 M⊙), the HZ remains entirely beyond the corotation radius at all ages.
However, for the 0.012 and 0.050 M⊙ brown dwarfs, the HZ migrates inward over time and intersects the corotation radius from Gyr (Figure 9). Any planet residing within what will become the HZ are expected to have already experienced strong tidal torques that drive inward orbital decay. As a result, planets may be removed through in spiral or disruption before the HZ completes its inward evolution, potentially leaving no planets remaining within the HZ at later times. Future studies must integrate both brown dwarf evolution and tidal migration with the modeling framework utilized in this work. This raises the question of which planets could actually have a lifetime in the HZ while residing within the corotation radius. Table 1 shows the calculated corotation radii for a select few brown dwarfs to provide reference. The corotation radius was calculated using Equation 3:
@@ IV Discussion
As shown in section III.1, previous studies such as Lingam et al. (2020) relied on analytic approximations rather than detailed brown dwarf evolution models, which did not accurately illustrate substellar HZ characteristics. Figure 1 illustrates this difference: at early ages, brown dwarfs are considerably cooler than the analytic models predict and thus much less likely to lose their entire water inventory than in the analytic case, although atmospheric escape timescales from brown dwarf planets requires further study. For the first time, equilibrium tracks for planets orbiting brown dwarfs are computed using state-of-the-art models that incorporate cloud physics and provide radius and luminosity evolution from 1 Myr onward. This permits us to explore the impact of clouds and deuterium burning on brown dwarf HZs. Since clouds slow the evolution of brown dwarfs, the cloudy evolution tracks result in orbiting planets that cool more slowly over time.
@@
Moreover, the transition from cloudy to cloudless states in the brown dwarf further delays planetary cooling at early ages, allowing well placed planets to maintain equilibrium temperatures in the HZ range for longer than predicted for cloudless evolution models. Lingam et al. (2020) concluded that planets orbiting brown dwarfs with masses less than 0.030 M⊙ would not maintain habitable temperatures over significant geologic timescales. This work illustrates that if cloud formation/dissipation, deuterium burning, and realistic atmospheric opacities are incorporated into atmospheric models, planets could remain in the HZ for hundreds of millions to billions of years (depending on the brown dwarf physical parameters, orbital radius, etc.), including planets orbiting brown dwarfs less than 0.030 M⊙ (shown in Figures 8 and 6).
@@
Sections III.2 and III.3 examined the duration that planets can remain within the HZ as a function of brown dwarf properties and orbital distance. Since higher-metallicity brown dwarfs cool more slowly and remain warmer at older ages than their lower-metallicity counterparts (e.g., Marley et al., 2021), planets orbiting metal-rich brown dwarfs spend longer periods within the HZ (see Figures 5 and 7). Consequently, planets orbiting metal-rich brown dwarfs spend longer periods within the HZ. This trend is evident in Figures 8 and 6, where systems with [M/H] = +0.5 exhibit extended habitable lifetimes compared to those with lower metallicities. The HZ lifetime also depends strongly on orbital distance and host mass: planets on wider orbits generally maintain habitability longer, particularly around lower-mass brown dwarfs. Figure 6 also shows the impact clouds have HZ lifetime. In the 0.001 au case shown in Figures 8 and 6, planets with masses above approximately 0.050 M⊙ exhibit longer lifetimes in Figure 8 than in Figure 6. This difference arises because the brown dwarfs at these masses are hotter (due to the formation of clouds), resulting in planetary temperatures exceeding 275 K.
@@
Section III.4 examines the evolution of the inner and outer HZ boundaries for both cloudless and cloudy brown dwarf models. In all cases, the HZ contracts and migrates inward as the brown dwarf cools over time. The inward migration and contraction of brown dwarf HZs shown in Figure 9 have important consequences when coupled with tidal evolution. For the 0.012 and 0.050 M brown dwarfs, the HZ intersects the corotation radius at late times, implying that planets residing in the HZ may experience strong inward tidal migration and be removed before long-term habitability can be sustained. In contrast, for the 0.080 M M dwarf, the HZ remains exterior to the corotation radius throughout its evolution.
@@
It is also important to note that connecting the hydrogen-burning limit of main-sequence stars to the brown dwarfs shifts the brown dwarf HZ farther outward, meaning the corresponding temperatures are higher. This occurs because Kopparapu et al. (2013) adopt lower albedos (and used a full radiative-convective model), which trigger a runaway greenhouse at lower levels of insolation and therefore place the inner edge of the HZ at smaller orbital distances. The planetary albedo vs. surface temperature of planets orbiting a variety of main-sequence stars is shown in Figure 6 in Kopparapu et al. (2013).
@@
In contrast, the shallower luminosity–mass relation of brown dwarfs produces nearly parallel HZ boundaries, yielding a relatively constant absolute width at younger ages. The systematically larger orbital distances of the main-sequence HZs are a direct consequence of their higher effective temperatures and bolometric luminosities, which shift the equilibrium temperature range for liquid water to wider separations. It is also clear that the HZ boundaries shift inward with time since brown dwarfs cool with age. Overall, brown dwarf HZs sweep through a large range of radii and Figure 10 illustrates how complex the evolution is.
@@
Although no rocky exoplanets have yet been discovered around brown dwarfs, studies have placed theoretical constraints on the occurrence rates of short-period planets in such systems (e.g., He et al. (2017). The transit method is particularly favorable for detecting rocky planets orbiting brown dwarfs: an Earth-sized planet transiting a brown dwarf with a radius of produces a transit depth of ppm, compared to only ppm for the same planet orbiting a G-type star. This large contrast underscores the strong potential for future searches targeting HZ planets around brown dwarfs. In future work, we will examine how the unique spectral energy distributions of brown dwarfs shape planetary climate by applying a 1D radiative–convective model to determine the inner (runaway greenhouse) and outer HZ limits for Earth-like planets (Windsor et al., 2022).
@@
However, the variability of the brown dwarf can cause the transit data to be extremely noisy. This is also seen in M-dwarf variability. There have been many studies analyzing the high activity and its effect on transits in M dwarfs (e.g., Yaptangco et al., 2025; Miyakawa et al., 2022). Similar studies to those listed could be a point of research for brown dwarfs as well. Considering that brown dwarfs have opposite evolution as main-sequence stars could provide another interesting window to the possibilities of habitability beyond Earth.
@@
Barnes and Heller (2013) highlighted fundamental challenges to habitability around cooling brown dwarfs (including tides). Our work further demonstrates that when cloud evolution and metallicity-dependent cooling are included, HZ lifetimes can be substantially longer than previously assumed. In contrast to Bolmont (2018), who emphasized that inward HZ migration coupled with tidal torques near the corotation radius may rapidly eliminate potentially habitable planets around brown dwarfs, our results demonstrate that when realistic brown dwarf evolutionary physics is taken into account, planets located beyond au, depending on the mass and corotation radius of the host brown dwarf, can both remain outside the corotation radius and persist within the HZ for extended durations (shown in Table 1).
@@ V Conclusions
We find that brown dwarfs are significantly less luminous at young ages than predicted by analytic approximations used in some recent brown dwarf HZ studies (Lingam et al., 2020). Those approximations neglect deuterium burning and cloud formation and dissipation, both of which play critical roles in determining accurate HZ boundaries and their evolution.
Deuterium burning and cloud formation and dissipation substantially extend HZ lifetimes for planets at specific orbital distances by maintaining those planets within the HZ while these processes operate.
Brown dwarfs with higher metallicities support longer HZ lifetimes because increased atmospheric opacity slows their cooling.
As brown dwarfs age, their HZs contract by orders of magnitude, depending on the brown dwarf mass.